# Spin Squeezing Enhanced Quantum Magnetometry with Nitrogen-Vacancy Center Qutrits

### v2

@@ II.2 Spin Squeezing Generation
To describe the intricate dynamics and capture the squeezing effects in the Hamiltonian for two NV centers, a squeezing term is introduced [36],
@@
where $c_{1}$ is the squeezing constant. The Hamiltonian, as depicted in Eq. (6), is referred as a one-axis twisting Hamiltonian. The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin. This, combined with the optimal initial state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
The generation of this type of Hamiltonian in real setup has been discussed in the literature. It can be realized through phonon mediation, as evidenced in [38] or produced in optical cavities [39]. In our specific protocol, we need to deactivate spin squeezing after a specific duration. A dynamic control method for NV center spin squeezing, as discussed in [40], could potentially address the technical challenge.
Alternative approaches to generating spin squeezing have been proposed. One example involves using microwave fields in silicon-vacancy centers, which are similar to NV centers [41]. Additionally, a universal method, introduced in [21], is also applicable to NV centers. The method presents a generic scheme for creating spin squeezing in any non-squeezing systems through the use of periodic rotation pulses.
@@ II.3 Dynamics
The NV center, while a robust quantum system, is inevitably influenced by the fluctuating states of nearby nuclear spins. These nuclear spins introduce a dynamic and unpredictable backdrop, contributing to the environmental noise that impacts the precision of measurements performed on the NV center. We assume very low temperatures relative to the NV center transition frequencies. Dephasing and thermalization are incorporated through the dissipators given as follows,
@@ III.2 Spin Squeezing Parameter
where $n_{\perp}$ is an axis perpendicular to the mean-spin direction $\hat{n}$, $N$ is the number of spins (in our case, $N=2$), and $\hat{J}$ represents the collective spin operator. Kitagawa-Ueda and other quantum correlations based criteria of spin squeezing [47] are commonly defined for SU(2) or spin-1/2 atomic ensembles. Spin squeezing in SU(3) or spin-1 systems can be discussed in their SU(2) subgroups, such as isospin squeezing [48]. We follow a similar approach to characterize quantum correlations and spin squeezing within the manifold of {|$\{\ket{0},\ket{-1}\}$. In general, spin squeezing can be optimized by rotating the coordinate frame and choosing the critical rotation angles. For simplicity, we only consider a fixed reference frame. Based on this manifold, the spin matrices $\hat{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ are written as
@@ IV.1 Analysis with spin-1 formalism
Furthermore, given the initial state is oriented along the $z$-axis, when spin squeezing is applied, it reduces the variance in a specific direction, creating an ellipsoid shape that remains aligned with the $z$-axis. A $\pi/2$ pulse then shifts the spin into the $x$-$y$ plane while preserving this variance ellipsoid. The effect of the magnetic field $B$ is to precess the spin around the $z$-axis, and the $x$-axis twisting deforms the noise cone, effectively narrowing the spin’s error ellipse. This sequence is illustrated in Fig. 4.

### v3

@@ II.2 Spin Squeezing Generation
To describe the intricate dynamics and capture the squeezing effects in the Hamiltonian for two NV centers, a squeezing term is introduced,
@@
where $c_{1}$ is the squeezing constant. The Hamiltonian, as depicted in Eq. (6), is referred as a one-axis twisting Hamiltonian [20, 36]. The choice of $x$-axis twisting in the context of spin squeezing is strategic, as it allows for the manipulation of the noise cone, effectively narrowing the error ellipse of the spin. This, combined with the optimal initial state, enhances the precision of the magnetic field measurement by interplaying with the $z$-axis rotation induced by the magnetic field [37].
II.2.1 Pulse Sequence Method
A more direct approach to generating spin squeezing is through pulse sequences that induce an axis-twisting effect. In [21] a method for achieving spin squeezing via such pulse sequences is described in detail. While their work focuses on spin-1/2 systems, this approach can be extended to spin-1 systems. This is particularly relevant to our setup, where we are interested in generating x-axis twisting. By adjusting the pulse properties, the method should allow us to transit from z-axis to x-axis twisting, which is more pertinent to the spin dynamics we are studying. A recalculation of the squeezing parameters within our parameter ranges will be necessary, but the method is sufficiently flexible to be implemented within our protocol.
II.2.2 Electric Field Induced Strain Method
An alternative method for generating spin squeezing, besides using pulse sequences, is the application of a strong electric field to induce strain in the system, leading to a Hamiltonian of the form $\sim[(S_{x})^{2}-(S_{y})^{2}]$. Although this Hamiltonian differs from the one used in our work, it still develops spin squeezing along specific axes, as demonstrated in [38].
II.2.3 Off-resonant Field Mediated Effective Interaction Method
Another approach, commonly used in spin-1/2 systems, involves effective interactions of the form $\sim S^{+}S^{-}+h.c.$, mediated by an off-resonant field that is adiabatically eliminated, as outlined in [39]. To generalize this to spin-1 systems, one can introduce strong laser fields and replace photon operators with c-numbers, yielding a Hamiltonian of the form $\sim g(S^{+})^{2}+h.c.$ Alternatively, large detuning can be employed via the Schrieffer-Wolff (SW) transformation, also known as the Fröhlich transformation, as shown in [40], to eliminate photon degrees of freedom, leading to a simpler effective Hamiltonian. This method has also been applied in systems like silicon-vacancy centers, which are similar to NV centers, as described in [41].
II.2.4 Other Methods
The generation of this type of Hamiltonian in real setup has been further discussed in the literature. It can be realized through phonon mediation, as evidenced in [42] or produced in optical cavities [43]. In our specific protocol, we need to deactivate spin squeezing after a specific duration. A dynamic control method for NV center spin squeezing, as discussed in [44], could potentially address the technical challenge.
Recent work [45] demonstrates the generation of large-scale spin squeezing from finite-temperature easy-plane magnetism, a method that shows great promise for NV centers.
To approximate a realistic measurement scenario, we calculate the Classical Fisher Information in Appendix B.
@@ II.3 Dynamics
To model the system, we adopt a Markovian noise framework and analyze the interplay between dephasing and thermalization coefficients in optimizing the squeezing effect. Our approach employs a phenomenological master equation incorporating decoherence and dephasing terms, consistent with standard treatments of bulk NV centers, which rely on jump operators linear in spin-1 defect operators ($S_{+}$, $S_{-}$, and $S_{z}$) [47, 48, 49]. The NV center, while a robust quantum system, is inevitably influenced by the fluctuating states of nearby nuclear spins. These nuclear spins introduce a dynamic and unpredictable backdrop, contributing to the environmental noise that impacts the precision of measurements performed on the NV center. We assume very low temperatures relative to the NV center transition frequencies. Dephasing and thermalization are incorporated through the dissipators given as follows,
@@
In this study, we focus on NV centers located deep within the bulk of the diamond, at depths exceeding tens of nanometers from the crystal surface. This assumption minimizes the influence of surface effects, such as charge noise caused by near-surface defects, allowing us to concentrate on bulk NV center dynamics. While charge noise is generally more pronounced for near-surface NV centers, its contribution decreases significantly with increasing depth, as demonstrated in [51]. This reference provides a detailed quantitative analysis of the depth-dependent impact of charge and spin noise on decoherence and relaxation rates. For instance, it shows that the electric field fluctuations due to surface charges scale inversely with the square of the defect depth. At depths greater than tens of nanometers, the noise contribution from surface charges becomes negligible compared to bulk noise sources. These findings support our assumption that the charge noise for deep NV centers is substantially less impactful. Additionally, experimental results in [52] confirm the depth dependence of relaxation and dephasing times, illustrating a marked improvement in coherence times as NV center depth increases. Figures 1 and 2 of [52] highlight this trend, with T2 and T1 coherence times improving by approximately two orders of magnitude as the depth increases from 4 nm to 12 nm.
@@ III.2 Spin Squeezing Parameter
where $n_{\perp}$ is an axis perpendicular to the mean-spin direction $\hat{n}$, $N$ is the number of spins (in our case, $N=2$), and $\hat{J}$ represents the collective spin operator.
In this work, we are using the particle-entanglement-based definition of spin squeezing, which is more restrictive than the noise redistribution version introduced by Wineland [56], commonly used in Ramsey measurements. The entanglement-based definition requires true multi-particle entanglement and cannot be achieved with a single spin, whereas the Wineland definition can reflect spin squeezing even in the absence of particle entanglement, merely by redistributing noise between different spin components.
Kitagawa-Ueda and other quantum correlations based criteria of spin squeezing [57] are commonly defined for SU(2) or spin-1/2 atomic ensembles. Spin squeezing in SU(3) or spin-1 systems can be discussed in their SU(2) subgroups, such as isospin squeezing [58]. We follow a similar approach to characterize quantum correlations and spin squeezing within the manifold of {|$\{\ket{0},\ket{-1}\}$. In general, spin squeezing can be optimized by rotating the coordinate frame and choosing the critical rotation angles. For simplicity, we only consider a fixed reference frame. Based on this manifold, the spin matrices $\hat{\sigma}=(\hat{\sigma}_{x},\hat{\sigma}_{y},\hat{\sigma}_{z})$ are written as
@@
Our approach highlights the quantum metrological advantage of employing higher-dimensional (spin-1) multipartite quantum entanglement. This advantage, distinct from conventional spin noise squeezing, is achieved by leveraging entanglement-based criteria for spin squeezing to further enhance sensitivity in quantum measurement protocols.
@@ IV.1 Analysis with spin-1 formalism
Our method is designed for measuring DC magnetic fields, where the typical readout technique is based on photoluminescence [26]. Additionally, optically detected magnetic resonance (ODMR) is commonly employed in such measurements [27, 61, 62].
@@
The spin system is initially oriented along the $z$-axis. The expectation values of the angular momentum operators are
$\langle J_{z}\rangle=\frac{N}{2},\quad\langle J_{x}\rangle=0,\quad\langle J_{y}\rangle=0,$ (22)
where $N$ is the total number of spins in the ensemble. In this state, the spin variance is symmetric around the $z$-axis, forming a circular uncertainty distribution in the $x$-$y$-plane. The uncertainty in each component is
$\langle(\Delta J_{x})^{2}\rangle=\langle(\Delta J_{y})^{2}\rangle=\frac{N}{4}.$ (23)
When spin squeezing is applied, the variance along one axis in the $x$-$y$ plane is reduced, while the variance along the orthogonal axis increases. The squeezed variances become,
$\langle(\Delta J_{\text{sq}})^{2}\rangle=\frac{N}{4\xi^{2}},\quad\langle(\Delta J_{\text{anti-sq}})^{2}\rangle=\frac{N\xi^{2}}{4},$ (24)
where $\xi^{2}<1$ is the squeezing parameter. The squeezed and anti-squeezed axes lie in the $x$-$y$ plane, but their exact orientation within the plane may not be specified. After applying a $\pi/2$ pulse, the spin state is rotated into the $x$-$y$ plane, and the noise ellipse is now aligned in this plane. The pulse does not alter the relative squeezing, but just the orientation of the state. The effect of the magnetic field $B$, applied along the $z$-axis, is to precess the spin state around the $z$-axis.
While we did not adopt a non-exponential decoherence model, such an approach could provide valuable insights by capturing slower, more realistic noise dynamics [63]. Furthermore, non-exponential decay could potentially extend coherence times, thus enhancing metrological precision in practical applications. Investigating this further may increase the relevance of our results to real-world quantum systems.
Moreover, we examine non-Markovian effects by introducing an ancillary qubit into the model, resulting in improved metrological performance, as discussed in Section IV.3.
@@ IV.2 Analysis with spin-1/2 formalism
IV.3 Non-Markovian Environment
To investigate non-Markovian effects in our system [64, 65], we introduce an ancillary qubit as a proof of principle. In our setup, this qubit is coupled to a single NV center through an $XX+YY$ interaction with a coupling strength of 0.1 GHz. The ancillary qubit, much like the NV center, is influenced by environmental effects such as dephasing and thermalization. However, in this case, the NV center experiences these environmental influences indirectly through its interaction with the ancillary spin.
By adding this single ancillary qubit, we effectively create a finite spin environment for the NV center. This finite environment can mimic a structured, non-Markovian bath, shifting the system’s response from a typical Markovian to a non-Markovian regime. As shown in Fig. 6, the QFI is significantly higher in the non-Markovian case, demonstrating enhanced sensitivity and suggesting that such structured environments may be advantageous for metrological applications.
In practice, NV centers are often surrounded by a finite number of nuclear or impurity spins, which can act as a non-Markovian environment and offer similar metrological benefits. By introducing an ancillary qubit as an effective finite environment, we provide a simplified model that captures these effects, highlighting the potential of non-Markovian dynamics for enhanced quantum sensing [66, 67].
Figure 6: For the spin-1 formalism, the red dotted curve shows the quantum Fisher information (QFI) as a function of time during the free evolution of a single NV center, including thermalization and dephasing with $\gamma_{t}=0.2$ MHz, $\gamma_{d}=0.1\gamma_{t}$, and $B=50$ G. The blue solid curve represents QFI under Markovian noise, while the red solid curve shows QFI under non-Markovian noise.
@@ V Conclusion
In light of surface effects, the dominant role of charge noise near the surface severely challenges the exploitation of spin squeezing and multipartite entanglement for magnetometry. Addressing these limitations requires exploring alternative strategies to enable quantum entanglement advantages for near-surface probes. Future research could focus on identifying optimal depths and surface properties to enhance magnetometry precision under non-Markovian dynamics and spin squeezing. These efforts are crucial for unlocking the full potential of NV centers for advanced sensing applications.
@@ Appendix A Entanglement Measure
Appendix B Classical Fisher Information
We calculate the Classical Fisher Information (CFI) during the system’s free evolution to better reflect a realistic experimental setup. Specifically, we evaluate the CFI using measurements along the eigenvectors of $S_{x}$ and $S_{y}$, which can be practically implemented through state tomography to reconstruct the quantum state and analyze its sensitivity to parameter changes. In quantum metrology, the CFI is a valuable metric for quantifying how responsive a measurement is to variations in a system parameter. For a probability distribution $\{p_{i}(\theta)\}$ that depends on a parameter $\theta$, the CFI is given by
Figure 8: The dark blue dashed line represents the Quantum Fisher Information (QFI) for a two-NV-center spin-1 system under spin squeezing conditions, with parameters $c_{1}=0.25$ GHz and $T_{\text{sq}}=30$ ns, accounting for thermalization and dephasing rates of $\gamma_{t}=0.2$ MHz and $\gamma_{d}=0.1\,\gamma_{t}$. The red dashed line indicates the QFI without spin squeezing. The cyan dotted line shows the Classical Fisher Information (CFI) using eigenvectors $S_{x}$ and $S_{y}$ as the POVM for a two-NV-center spin-1 system with spin squeezing (parameters: $c_{1}=0.25$ GHz, $T_{\text{sq}}=30$ ns), while the solid magenta line represents the CFI without spin squeezing.
$\text{CFI}(\theta)=\sum_{i}\frac{1}{p_{i}(\theta)}\left(\frac{\partial p_{i}(\theta)}{\partial\theta}\right)^{2},$ (31)
where $p_{i}(\theta)$ represents the probability of the $i$-th measurement outcome for a given $\theta$. In our context, these probabilities correspond to outcomes measured along the eigenvectors of $S_{x}$ and $S_{y}$, providing the necessary basis for CFI calculation. By choosing eigenvectors of $S_{x}$ and $S_{y}$ as the Positive Operator-Valued Measure (POVM) elements, we model measurements that are feasible in practical setups. Such measurements can be carried out using state tomography, as detailed in [72]. In Fig. 8, we observe that the CFI is non-zero and is higher with squeezing, which further supports our results.